OpenStudy (anonymous):

I walked East 5 meters and turned Northward to reach Durham. Which was 10 meters. What angle did I make to turn to go to Durham? Use Cos-1 or inverse cosine to get your answer.

OpenStudy (anonymous):

In order to use the cosine of the angle, you need the hypotenuse of the triangle. If East is the x direction, and North is the y direction, use X^2 + y^2 to find the square of the hypotenuse. Let me know if you need more help from there.

OpenStudy (anonymous):

yes please I figured out that the side is 8 but I dont know how to use the cos-1 functions on the calculator can u help me out please thank you

OpenStudy (anonymous):

Cos(theta) = h/x theta = Cos-1(h/x) So divide h (the hypotenuse) by x, then find on your calculator the Cos with a -1 and just push it.

OpenStudy (anonymous):

sorry .. it's x/h Cos(theta) = x/h

OpenStudy (anonymous):

thanks ill try it out

OpenStudy (anonymous):

The hypotenuse doesn't equal 8. The side is: \[ \sqrt{5^2 + 10^2} = \sqrt{125} = 11.2\]

OpenStudy (anonymous):

ok thank you...im not very good at this :(

OpenStudy (anonymous):

Did you find the angle?

OpenStudy (anonymous):

no im giving up :/ this is too hard for me

OpenStudy (anonymous):

Wait.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Divide 5 by 11.2 Then push the cos with a -1 button on your calculator.

OpenStudy (anonymous):

You might have to press the 'shift' key on your calculator, then press Cos, if Cos with a -1 is written above the Cos button.

OpenStudy (anonymous):

when I do that it sayd invalid imput

OpenStudy (anonymous):

ok i ithink the answer is 63.49

OpenStudy (anonymous):

That's correct! Congratulations.

OpenStudy (anonymous):

when I round it off

OpenStudy (anonymous):

oh wow your amazing!!

OpenStudy (anonymous):

aww thnx so very much for helping me...I was going to give up but you didnt let me thnxx:) !!!!

OpenStudy (anonymous):

You're Welcome!